how do i convert this linear equation into slope-intercept form?
2x-3y=7


Thank you!!

Answer by Alan3354(69443)

how do i convert this linear equation into slope-intercept form?
2x-3y=7
-------------
You solve for y. That's all that's ever necessary.
2x-3y=7
-3y = -2x + 7
y = (2/3)x - 7/3 (y = mx + b)
That's the whole secret.